OpenStudy (anonymous):

Solve. x^2 + 8x + 25 = 0

hero (hero):

I'm pretty sure this is going to end up being an imaginary solution

OpenStudy (campbell_st):

looks like it 4 +- 3i

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^{}-4ac}}{2a}\] why cant v use this..

OpenStudy (anonymous):

offcourse you can buddy

OpenStudy (anonymous):

nope i dont think...maybe the equation is wrong...

OpenStudy (anonymous):

why not you can

OpenStudy (anonymous):

try and c :)

OpenStudy (campbell_st):

just you'll get a negative radical... so you'll need complex number notation

OpenStudy (anonymous):

\[(-8+-\sqrt{64-4\times25})\div2\times1\]

OpenStudy (anonymous):

may you are not familiar with complex numbers

OpenStudy (campbell_st):

then the solution is undefined

OpenStudy (anonymous):

it's not a great deal

OpenStudy (anonymous):

square root of negative numbers can be evaluated by defining anew number called IOTA denoted by i (refer the answer given by campell you can see an i there ) this i is nothing but square root of -1, \[\sqrt{-1}=i\] so you can modify \[\sqrt{-9}=\sqrt{9}\times \sqrt{-1}\] but we already know that \[\sqrt{-1}=i\] so \[\sqrt{-9}=\sqrt{9}i=3i\]